2ln square root x-8=8
Answer by Boreal(15235)

2 ln *sqrt(x-8)=8
ln*sqrt(x-8)^2=8
raise both sides to e power
sqrt(x-8)^2=e^8;
x-8=e^8
x=8+e^8=2988.96
=========
check
2 ln sqrt (2980.96)=8
the sqrt term is 54.6
the ln of the sqrt term is 4
2*4=8
Answer is 8+e^8